Two runners leave from the same location traveling in opposite directions. The first runner is running at 8 mph and the second is running at 10 mph. How long have they been running when they are 36 miles apart?
Answer by mananth(12270)

They are running in opposite directions
so effective speed = 10+8 = 18 mph
d=36
time = 36/18 = 2 hours